OpenStudy (littlet):

Find the equation of a line perpendicular to Y=5x+3. that passes through the point (-15,8)

OpenStudy (mathmale):

Glad to have you as a new member of OpenStudy, LittleT! Welcome. Given the line y=5x+3, find the slope of this line. call it m1. To find the slope of a line perpendicular to the given line, take the negative reciprocal of m1. Have you done that before? Call the result m2. m2 = ?

OpenStudy (littlet):

Thank you and this is the perpendicular proof right?

OpenStudy (mathmale):

Not a proof, no. We're finding the slope of the given line, and then using it to find the slope of a line perpendicular to the given line. Have you been able to identify m1?

OpenStudy (mathmale):

@littlet?

OpenStudy (littlet):

Ok sorry. is M1 not y=5x+3?

OpenStudy (mathmale):

See my private message. m1 is 5 alone. y=mx+b, where m=slope and b=y-intercept.

OpenStudy (littlet):

is m2 -5?

OpenStudy (littlet):

And yes i did see your message.

OpenStudy (mathmale):

Given: m1=5 m2 is the "negative reciprocal" of m1. This means you must first invert m1 to get 1 / m1, and then change the sign. Start with m1=5. Invert 5. Result?

OpenStudy (littlet):

-1/5

OpenStudy (mathmale):

That's right. That's your m2. Now you have the slope of a line perpendicular to the given line. You also have a point thru which the perpendicular line passes. m2 = -1/5, and the point is (-15,8)

OpenStudy (mathmale):

Pleae use the point-slope formula for the equation of a straight line to write the equation of the perpendicular line.

OpenStudy (mathmale):

The point-slope formula is \[y-y _{0}=m(x-x _{_{0}})\]

OpenStudy (mathmale):

Here your x0 = -15 and your y0=8

OpenStudy (littlet):

OK so x0 = -15 and your y0=8 is the equation in point-slope formula?

OpenStudy (mathmale):

Not quite, LT. Start with \[y-y _{0}=m(x-x _{_{0}})\]

OpenStudy (mathmale):

and substitute -15 for x0, and 8 for y0. Try it, please.

OpenStudy (littlet):

y-8=m(x- 15)

OpenStudy (littlet):

that right?

OpenStudy (mathmale):

\[y-( ? ) = m*(x -( ? ) )\]

OpenStudy (mathmale):

almost right. remember that x0 is -15. Have you taken that into account?

OpenStudy (mathmale):

What is x-(-15)?

OpenStudy (littlet):

Yes sorry i ment -15

OpenStudy (mathmale):

so you should end up with x+15. x-(-15)=x+15.

OpenStudy (mathmale):

So now you have \[y-8=m(x+15)\]

OpenStudy (littlet):

ok

OpenStudy (mathmale):

Pleae go back and find our value for m2. Subst. that into the equation I've just shared with you.

OpenStudy (littlet):

m2 was -1/5 where would that go?

OpenStudy (mathmale):

\[y-y _{0}=m2(x+15)\]

OpenStudy (littlet):

oh ok so y- 8 = -1/5(x+15)

OpenStudy (mathmale):

got to memorize that equation (point-slope form):\[y-y _{0}=m(x-x _{0})\]

OpenStudy (mathmale):

Your result is perfect except that that -1/5 has to go inside parentheses for clarity.

OpenStudy (mathmale):

\[so: y- 8 = -(1/5)(x+15)\]

OpenStudy (littlet):

alright got it. so tat is the solution?

OpenStudy (mathmale):

This line goes thru the point (-15,8) and has a slope -1/5, which is the negative reciprocal of the slope of the given line.

OpenStudy (mathmale):

Yes, that's the solution.

OpenStudy (mathmale):

;)

OpenStudy (littlet):

Thank you so much!! have a great night!

OpenStudy (mathmale):

You're welcome in more ways than one. Good night to you too (but it's only 3:23 p.m. here in California, where I live). See you again on OpenStudy!

OpenStudy (littlet):

Oh lol alright. Cya and thank you again .

OpenStudy (mathmale):

take care! Bye!